OpenStudy (anonymous):

How to rewrite the following as f(x)=mx+b x -y - 5 =0

OpenStudy (anonymous):

you want y to be in the place of f(x). f(x) can basically be seen as a way denote a complex version of y (don't take time to memorise that, haha!). basically, when f(x)=some random number; then y = some random number so. y=f(x) i.e f(x) = mx + c y = mx + c so taking the equation y - x - 5 = 0 -y = -x + 5 y = x - 5 Remember that m of mx is the gradient of the line (so it can be equal to 1 in which case we don't write down m) and c is some constant.